OpenStudy (anonymous):

If y − 4 = 2x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(2, 1), (4, 2), (6, 3)} {(1, 2), (2, 4), (3, 6)} {(4, 0), (6, 1), (8, 2)} {(0, 4), (1, 6), (2, 8)}

OpenStudy (anonymous):

Plug them into the equation and see if they work///

OpenStudy (anonymous):

All of the outputs for each functions would be in order 0, -2, 4, and -4 @nelsonjedi

OpenStudy (anonymous):

y is the answer for 2x and subtract it by 4? I think I'm doing it wrong

OpenStudy (anonymous):

Look at the first group. Take the first ordered pair (2,1) in plug into y-4=2x....or 1-4 = 2(2)..is this a true statement?

OpenStudy (anonymous):

I don't think so, 1-4 would be 3 and 2x2 would be 4?

OpenStudy (anonymous):

So it'd be D? If I plug in D it'd be 4-4=2x0 or 0=0

OpenStudy (anonymous):

Do all the pairs in D work?

OpenStudy (anonymous):

Yes, (0,4)=0, (1,6)=2, and (2,8) would be 4.

OpenStudy (anonymous):

Bingo..good job

OpenStudy (anonymous):

Thank you, It seems my mistake was that I didn't know that you plug in the Y, because I'm used to the y being the answer.